OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of 2x2 + 4x + 3 = 0.

OpenStudy (anonymous):

for a quadratic equation of the form ax^2+bx+c=0 the discriminant is b^2-4ac

OpenStudy (anonymous):

If \(\large b^2-4ac \) is: positive, then you'll two real solutions; zero, one real solution; negative, then you'll have two complex solutions.

OpenStudy (anonymous):

ya dpalnc is right

OpenStudy (anonymous):

Quadradic: \[\huge x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] It's the part @nitz and @dpaInc said, with the deductions you'll be using that @dpaInc said. :-)

OpenStudy (anonymous):

And technically if it's zero, it's one double-root solution (the same solution twice). But for most practical purposes you don't have to worry about that.

OpenStudy (anonymous):

but agent x she has to find the nature of roots not the roots itself